HH and נR are fashion designers who agreed to form a partnership to open a dothing store. An attorney prepares the partnershi Answer all
Answer #1

Answer-3

HH JR

Cash A/c Dr.

Accounts receivable A/c Dr.

Allowance for uncollectible accounts (FV) Dr.

Supplies Dr.

Equipments (FV) Dr.

To Accounts Payable

To Capital Account

80000

104000

6000

2000

40000

64000

168000

Cash A/c Dr.

Accounts receivable A/c Dr.

Allowance for uncollectible accounts (FV) Dr.

Supplies Dr.

Equipments (FV) Dr.

To Accounts Payable

To Capital Account

60000

40000

4000

1000

20000

18000

107000

Narration for both : Being Assets recorded at FV and Liability recorded at Book Value

Answer - 4

Total Captial after admission = 200540 + 145460 + 50000 = 3,96,000

a) Purchase from Partners :

Paid to HH (50000 * 60%) = 30000

Paid to JR 50000 * 40% = 20000

Entries :

HH's Captial Ac Dr.

Goodwill

To LV

20054

9946

30000

(being 10% purchased from HH)

JR's Captial Ac Dr.

Goodwill

To LV

14546

5454

20000

(being 10% purchased from JR)

b)

HH's Capital Ac Dr.

JR's Captial Ac Dr.

Goodwill

To LV's Capital

20054

14546

15400

50000

((being LV admitted to partnership)

c) Total capital after admission = 200540+145460+50000 = 396000

15% of 396000 = 59400

HH's Capital Ac Dr.

JR's Captial Ac Dr.

Cash

To LV's Capital

5640

3760

50000

59400

(Being LV admitted for 15% and loss beared by HH & JR)

d) Total Capital before admission = 346000

60% of 346000 = 207600

Amount received from LV = 350000

Goodwill = 350000-207600 = 142400

HH's Capital Ac Dr.

Goodwill (HH's Share) Dr.

Goodwill (207600-200540) Dr.

To LV's Capital

200540

142400

7060

350000

(being HH share purchased by LV)

Answer - 5

(1)

Thomas's Capital A/c Dr.

To Ruby

To Stanley

To Goodwill

20000

9000

9000

2000

(2)

Thomas's Capital A/c Dr.

To Cash Ac

To Goodwill

20000

16000

4000

(3)

Thomas's Capital A/c Dr.

Goodwill Dr.

To Cash

20000

5000

25000

Answer - 6

Amount Received by Tom = 100000

Capital of Tom = 60000

Goodwill received = 100000-60000 = 40000

Tom's Share in firm = 25%

Tom Received 40000 for 25%

Therefore total Goodwill of firm = 40000/25% = 160000

Total Goodwill = $ 1,60,000